A basketball player practices shooting three-pointers from a distance of 7.50 m from the hoop, releasing the ball at a height of 2.00 m above the floor. A standard basketball hoop’s rim top is 3.05 m above the floor. Th e player shoots the ball at an angle of 48.0° with the horizontal. At what initial speed must she shoot to make the basket?


1
Expert's answer
2021-12-27T08:00:49-0500

"\alpha = 48\degree"

"h_0 = 2m"

"h_1 =3.05m"

"g = 9.8\frac{m}{s^2}"

"s_1= 7.5"

"\vec v-?"

"s = v_x*t= v*\cos \alpha* t"

"t= \frac{s} { v*\cos \alpha}"

"h_1 = h_0+v_yt-\frac{gt^2}{2}"

"v_y = v*\sin \alpha"

"h_1 = h_0+s*\tan \alpha-\frac{gs^2}{2*v^2*\cos^2\alpha}"


"3.05 = 2 +7.5*1.11-\frac{9.8*7.5^2}{2*v^2*0.67^2}"

"v = 9.23\frac{m}{s}"


"\text{Answer: }9.23\frac{m}{s}"